# augmented reality AP average precision

### v1

@@ III-C Gaussian Rasterization
Our rasterization process consists of two key stages. First, we build upon semantic GS. This initialization and pre-training phase enables the model to distinguish and categorize scene elements effectively, leveraging semantic information for a strong foundation in subsequent refinements.
In the main phase of our approach, which focuses on PoI refinement, we utilize the standard GS rasterizer while employing optimizations to enhance efficiency. Specifically, we accelerate the training process using a fused-SSIM loss function alongside a sparse Adam optimizer, significantly improving convergence speed [14]. Additionally, we integrate our PoI masking mechanism, allowing for targeted refinement of relevant regions while optimizing a clean scene representation and loss computation. This selective approach ensures that our approach concentrates on relevant areas.
IV EVALUATION
IV-A Metrics
To evaluate the rendering quality of our approach, we use the metrics peak signal-to-noise ratio (PSNR) and similarity index measure (SSIM[39]. For the update time, we specify the total training time in ($s$) and iterations per second ($it/s$). In Table I and Table II, differences in PSNR are marked as: $\pm, 1$.0-1.0, and $\leq 1.0$. For time in seconds we mark the lowest values within $\pm$30 seconds, values larger but below $\geq 100$ seconds, >100>100 and <1000<1000 seconds, and values $1000>$ seconds. The best values per metric are marked in bold.
IV-B Implementation Details
Hardware Setup
Our approach and the comparison methods (pure GS [14] and $G$ [15]) are implemented in Python with PyTorch and CUDA. All scenes were evaluated on desktop PCs with an RTX3090 with 24 GB VRAM.
Periodic Scene Filtering
We use parallelization to calculate $p^{*}$. For this we use a thread pool with $N$ workers so that several views can be processed simultaneously. To maintain visual fidelity during the refinement process, periodic filtering of the Gaussian based on color similarity is performed every 1000 training iterations.
TABLE I: Evaluation on NeRDS 360 [20]. We report PSNR, SSIM, and Time in seconds. We compare $G$ with a convex-hull removal and $G$ with a direct removal scheme with our method.
Scene GG [15] (convex-hull) GG [15] (direct removal) Ours
[…]
6thAndMission_medium
6 11.403 0.724 2560 16.662 0.828 2560 26.430 0.962 418
7 11.403 0.724 1957 21.726 0.904 1957 28.106 0.968 469
10 11.117 0.688 2061 22.378 0.896 2061 27.466 0.964 447
GrantAndCalifornia
1 9.729 0.753 1848 20.036 0.885 1848 25.989 0.959 446
2 9.479 0.703 2005 19.371 0.879 2005 26.876 0.959 446
3 9.701 0.744 1867 18.096 0.853 1867 28.053 0.965 443
VanNessAveAndTurkSt
3 12.276 0.788 1939 17.802 0.852 1939 26.197 0.957 464
5 11.097 0.753 2160 19.040 0.865 2160 28.518 0.974 480
Mean 11.238 0.748 2050 19.389 0.870 2050 27.204 0.963 460
Figure 3: Exemplary occlusion removal comparison on the scenes SF_GrantAndCalifornia1 (bottom) SF_6thAndMission_medium10 (top) after 30,000 iterations. Post-training isolation using Gaussian Grouping (b, c) partially recovers occlusions, with convex hull usage influencing the result.
@@ IV-B Implementation Details
Figure 4: Initialization example on SF_6thAndMission_medium10 from the Neo360 dataset [20]. The rendering iteration is denoted among each column.
@@
Figure 5: SCRREAM scene01 after 30k iterations (images cropped to PoI). GG using a convex hull (a), and GG using a non-concex hull (b), and our PoI refinement (c).
IV-C Datasets
NeRDS 360 [20]
We use the COLMAP subset from the NeRDS 360 dataset [20]. The scenes are particularly interesting as they show urban landscapes and vehicles that resemble the conditions of possible rescue situations and offer a 360 view of a street scene. The respective scenes with COLMAP ground truth are two from SF_VanNessAveAndTurkSt (3,5), the scenes 1, 2, and 3 from SF_GrantAndCalifornia, and three scenes from 6th and Mission (6, 7, 10). As PoI we defined the class “car”.
SCRREAM [19]
For an indoor proof of concept, we evaluated our approach on SCRREAM [19]. Each scene contains separate recordings with a variety of objects. For our comparison we use Scene01, Scene02, and Scene03 all in the setting full. We use the class chair as PoI.
IV-D Baselines
For the comparison, we selected $G$ [15], as it is the state-of-the-art in GS separation and showed the same performance in NVS as traditional GS.
IV-E Novel View Synthesis Quality
We compare the overall performance on all datasets with a focus on our speedup compared to semantic GS, evaluate the used initialization iterations using both baselines, and compare our approach to traditional GS in terms of NVS and runtime.
TABLE II: Evaluation on SCRREAM [19]. For all scenes, we use the full configuration (_full_00). We report PSNR, SSIM, and Time in seconds. As a comparison method, we choose $G$ with a direct object removal.
Scene GG [15] Ours
[…]
scene01 31.90 0.927 2614 30.89 0.903 709
scene02 25.61 0.849 2766 25.97 0.792 1083
scene03 36.52 0.965 2645 35.79 0.954 656
Mean 31.34 0.914 2675 30.88 0.883 816
TABLE III: Comparison of our approach with different initialization iterations. We evaluated after 7k and 30k iterations on scene SF_6thAndMission_medium10 [20]. We marked the best approach bold. We report PSNR, SSIM and Time in seconds.
initial 7k iterations 30k iterations
@@ IV-E Novel View Synthesis Quality
IV-E1 Comparison to Full Semantic Gaussian Splatting
Since our method leverages semantic GS for initialization, we compare it with the respective method, using the complete training cycle.
With regard to the quality of NVS, we focus on PSNR and SSIM, where PoI masking can be applied. This ensures a fair comparison when calculating these metrics. As shown in Table I, our approach outperforms GG in three scenes while achieving competitive results overall in terms of SSIM. For PSNR, both methods show similar ranges, approximately $\pm 0.5$. However, our approach not only achieves this level of performance but also demonstrates a significantly reduced training time, particularly excelling in performance for the “car” class.
To evaluate our approach in multiple settings, we also compared it in an indoor dataset. The SCRREAM dataset [19] (real-world, indoor) consists of diverse scenes with highly accurate ground truth for semantic segmentation. As shown in Table II, our approach outperforms GG in PSNR on one scene and remains competitive across all three scenes in both PSNR and SSIM. As shown in Figure 5, no visual downside is visible for our approach; rather, we show clearer boundaries.
In Figure 3, we provide a direct comparison of filtered outputs between GG and our method. A comparison of our PoI approach with GG using convex hull removal reveals that the convex hull method retains numerous outliers (floater) from the original scene. Similarly, when using non-convex, class-based removal, additional outliers become evident, as shown in Figure 3. Despite these differences, our method exhibits no significant drawbacks in terms of NVS quality.
IV-E2 Impact of Varying Numbers Initialization Iterations
The visual quality shows that using a low-quality semantic GS scene for initialization can reduce the final quality within a range of $\pm 0.5$. Although this difference is per se not high, when looking at Figure 4 for a visual scene exploration, more iterations are beneficial. 3000 to 4000 iterations show clearer results. Thus, 3000 iterations provide a positive balance between training time and PSNR and SSIM result.
TABLE IV: Comparison of GS after 15k and 30k training iterations. We mask PSNR and SSIM with the GT object mask of the car, as GS does not offer per-object extraction. We report the best result in bold.
Iterations PSNR $\uparrow$ SSIM $\uparrow$ TIME [s] $\downarrow$
GS 30000 21.40 0.862 417
Ours 15000 21.73 0.859 285
IV-E3 Comparison to Traditional Gaussian Splatting
Traditional GS [14] has seen significant improvements, including a reduction in training time. However, while segmentation-based removal is not possible in traditional GS, it is possible to mask the metric calculation for a comparison in quality. When comparing this method with our approach, see Table IV, we demonstrate that our hybrid scene representation, initialized with semantic GS, outperforms traditional GS in both runtime and PSNR. For this comparison, we used masked PSNR and masked SSIM as traditional GS does not allow the removal of specific scene objects. Our approach achieves a better PSNR in just half the number of training iterations, while SSIM is almost the same performance.
IV-F Speedup Comparison
Adding segmentation to the GS rasterization process can enhance quality, but also increases the training time. As shown in Table I, GG around 2050 seconds to process all scenes compared to our approach, which requires 460 seconds. Our approach only requires a segmentation and a ready-to-explore scene (3k), and we achieve a maximum training time of around 460 seconds, where around 120 seconds can be attributed to the initialization training and around 4 seconds to the PoI extraction. This shows that our approach reduced the training load enormously.
@@ V DISCUSSION
We propose a novel approach for PoI in GS, leveraging semantic GS [15] in the first stage to initialize the PoI refinement step. The semantic GS module is interchangeable and independent from the PoI refinement, fostering modularity and extendability of the approach. Our masking and PoI selection allow for an efficient training cycle. The PoI selection ensures that only relevant data for the target object or target area contributes to the scene optimization, which allows a more focused and faster training.
As shown in Table I, the runtime of GG is notably higher compared to our approach. While semantic information is necessary for selecting one PoI or multiple PoIs after obtaining the rough scene representation, our method provides a substantial speedup, performing faster than training full GG and extracting more visually appealing PoIs (Figure 3).
Experiments on the SCRREAM dataset validate the robustness of our method in different environments. While its relative advantage over the baselines is slightly smaller compared to the outdoor NeRDS 360 scenes, our method still produces strong results.
Moreover, the PoI refinement gives an advantage for occluded areas, while in the full scene, a pole or street light could occlude the PoI from different NVS perspectives. Selecting a PoI and extracting it shows the advantage of our approach against occlusion. This capability offers a more complete understanding of the object’s geometry and structure from the same perspective, a feature not achievable with GG.
@@ VI LIMITATIONS
Our approach relies on high-quality segmentation, and its effectiveness may be constrained when segmentation accuracy is insufficient. Previous work [4] has already demonstrated the critical role of segmentation labels in GS-based methods, highlighting their impact on reconstruction quality and downstream tasks.
@@ VII CONCLUSION
In summary, we introduce a PoI refinement framework for GS that selectively enhances segmentation-based regions critical for scene exploration. By focusing computation on targeted scene areas, our approach produces higher-fidelity, PoI without the overhead of refining the entire scene representation. Our evaluation on two benchmarks demonstrates clear advantages over both standard GS and semantic GS. Unlike standard GS, which lacks PoI support, our method achieves targeted visual refinement, and compared to semantic GS, it reduces training time while preserving segmentation accuracy and rendering quality. These results highlight our approach’s efficiency and visual quality.

### v2

@@ III-C Gaussian Rasterization
Our rasterization process consists of two key stages, see Fig. 2. First, we build upon semantic GS, in our case GG [8]. This initialization and pre-training phase enables the model to distinguish and categorize scene elements effectively, leveraging semantic information for a strong foundation in subsequent refinements.
In the main phase of our approach, which focuses on POI refinement, we utilize the standard GS rasterizer while employing optimizations to enhance efficiency. Specifically, we accelerate the training process using a fused-SSIM loss alongside sparse Adam optimizer, significantly improving convergence speed [7]. Additionally, we integrate our POI masking mechanism, see Section III-A, allowing for targeted refinement of relevant regions while optimizing a clean scene representation and loss computation. This selective strategy ensures that our approach concentrates on relevant areas.
@@
Figure 4: Initialization example on SF_6thAndMission_medium10 from the Neo360 dataset [14]. The rendering iteration is denoted below each column.
TABLE I: Evaluation on NeRDS 360 [14]. We select “car” as POI.
GG [8] (“convex hull”) GG [8] (direct removal) Ours
[…]
6thAndMission_medium
6 12.545 0.547 0.331 2560 20.184 0.724 0.245 2560 27.776 0.959 0.061 418
7 18.858 0.753 0.187 1957 25.732 0.828 0.139 1957 29.790 0.967 0.045 469
10 10.808 0.484 0.366 2061 23.164 0.865 0.162 2061 27.466 0.964 0.058 447
GrantAndCalifornia
1 11.324 0.433 0.298 1848 24.603 0.821 0.164 1848 32.094 0.968 0.052 446
2 10.890 0.433 0.354 2005 22.567 0.758 0.184 2005 31.292 0.960 0.058 446
3 10.341 0.445 0.349 1867 20.030 0.711 0.221 1867 29.822 0.963 0.051 443
VanNessAveAndTurkSt
3 13.459 0.601 0.275 1939 24.469 0.750 0.197 1939 29.010 0.957 0.069 464
5 17.131 0.602 0.271 2160 24.469 0.750 0.197 2160 29.269 0.971 0.047 480
Mean 13.170 0.537 0.304 2050 22.636 0.773 0.191 2050 29.590 0.964 0.055 460
IV EVALUATION
IV-A Metrics
To evaluate the novel view synthesis (NVS) quality of our approach, we report peak signal-to-noise ratio (PSNR), learned perceptual image patch similarity (LPIPS), and similarity index measure (SSIM). We report the total training time in seconds.
IV-B Implementation Details
All scenes were evaluated on desktop PCs with an RTX3090 with 24 GB VRAM.
For our periodic scene filtering, we use parallelization to calculate $p^{*}$. For this, we use a thread pool with $N$ workers so that several views can be processed simultaneously. To maintain visual fidelity during the refinement process, periodic filtering of the Gaussians based on color similarity is performed every 1000 training iterations.
IV-C Datasets
NeRDS 360 [14]
We use the COLMAP subset from the NeRDS 360 dataset [14]. The scenes are particularly interesting as they show urban landscapes and vehicles, while offering a 360 view of a street scene. The respective scenes with COLMAP ground truth are two from SF_VanNessAveAndTurkSt (3,5), the scenes 1, 2, and 3 from SF_GrantAndCalifornia, and three scenes from 6th and Mission (6, 7, 10). As POI we defined the class “car”.
SCRREAM [13]
For an indoor proof of concept, we evaluated our approach on scenes from the SCRREAM [13]. Each scene contains separate recordings with a variety of objects. For our comparison we use Scene01, Scene02, and Scene03 all in the setting full. We use the class chair as POI.
IV-D Baselines
For the comparison, we selected GG [8], as it is the state-of-the-art in GS separation. Furthermore, as reported by Ye et al.[8], it has the same performance in NVS as vanilla GS. We exclude other approaches like GAGA [12], due to several needed subsequent training steps, and simultaneous localization and mapping (SLAM)-based approaches [33], as they require longer training times.
TABLE II: Comparison of GS after 15k and 30k training iterations. We mask PSNR and SSIM with the GT object mask of the “car”, as GS does not offer per-object extraction.
Iterations PSNR $\uparrow$ SSIM $\uparrow$ TIME $\downarrow$
GS 30000 21.40 0.862 417
Ours 15000 21.73 0.859 285
IV-E Runtime
IV-E1 NeRDS 360
To evaluate the runtime, we measure the overall training time of our approach (CoRe-GS), and the compared baseline (GG with “convex hull” removal, and direct removal), including time for semantic object extraction. As shown in Table I our overall runtime is 460 seconds, while GG, requires 2050 seconds in total. Adding segmentation to the GS rasterization process can enhance quality, but also increases the training time. Our approach only requires a segmentation and a ready-to-explore scene (3k), and we achieve a maximum training time of around 460 seconds in average, where around 120 seconds can be attributed to the initialization training and around 4 seconds to the POI extraction. This shows that our approach reduced the training load enormously.
IV-E2 SCRREAM
Similarly, on the SCRREAM dataset, see Table IV, we require 816 seconds on average while GG takes 2675 seconds. On the SCRREAM dataset, our runtime is more influenced by the number of images used for the POI refinement compared to the NeRDS 360 dataset. The number of images per scene, and per POI is mostly similar on NeRDS 360, while for the non-360 scenes (SCRREAM) the images containing POIs vary more.
(a) No filtering
(b) Filtering
Figure 5: Color-based filtering (cropped to outlier) on SF_GrantAndCalifornia2. Our approach without filtering shows outliers (left), while our filtering leads to a clean background (right).
IV-E3 Comparison to Vanilla Gaussian Splatting
GG and in general semantic GS have prolonged training times due to the extended rasterization process. Since its introduction, the codebase of vanilla GS [7] has seen significant improvements, including a reduction in training time. When comparing this vanilla GS with our approach (CoRe-GS), see Table II, we demonstrate that CoRe-GS, initialized with semantic GS, outperforms vanilla GS in both runtime and PSNR. For this comparison, we used masked PSNR and masked SSIM as traditional GS does not allow the removal of specific scene objects. Our approach achieves a higher PSNR in just half the number of training iterations, while SSIM shows almost the same performance. Furthermore, the overall training time is lower when using our approach.
IV-F Optimizations
CoRe-GS proposes several optimizations for semantic object extraction. While we show the impact on runtime and NVS quality separately on two datasets, we also focused on the joint impact of these. In the following, we evaluate the impact of our chosen number of initializations, as well as the impact of our proposed color-based filtering.
IV-F1 Initialization
The visual quality shows that using a low-quality semantic GS scene for initialization can reduce the final quality within a range of $\pm 0.5$ for PSNR. Although this difference is not high per se, when looking at Figure 4 for a visual scene exploration to select POIs, the benefits of additional iterations become clear. 3000 to 4000 iterations show clearer results. Similarly, in metrics, we can observe that when even using 3000 iterations, combined with our proposed optimizations, a better NVS quality can be achieved. Thus, 3000 iterations provide a positive balance between training time and NVS quality.
IV-F2 Filtering Impact
In Fig. 5, we observe noticeable outliers when color-based filtering is not applied, compared to when it is used. Quantitatively, for the scene SF_GrandAndCalifornia2, the results without filtering are 31.038 PSNR, 0.959 SSIM, and 0.060 LPIPS. With our color-based filtering, PSNR improves by 0.253, SSIM improves by 0.001, and LPIPS improves by 0.002. Although the metric gains are most pronounced for PSNR, the qualitative improvement in Fig. 5 clearly highlights the reduction of visible outliers.
IV-G Novel View Synthesis Quality
IV-G1 NeRDS 360
We compare the overall performance on all datasets with a focus on our speedup compared to GG. As shown in Table I, our approach outperforms GG. When using both settings for GG (“convex hull”, and direct removal). Our approach not only achieves this level of performance but also demonstrates a significantly reduced training time. We outperform GG in all visual metrics. In Figure 3, we provide a direct comparison of filtered outputs between GG and our method. A comparison of CoRe-GS with GG using “convex hull” removal reveals that the “convex hull” method retains numerous outliers (floater) from the original scene. Despite these differences, our approach (CoRe-GS) exhibits no significant drawbacks in terms of NVS quality. The “convex hull” clearly leaves in scene parts, one would like to remove, like the street. While the “direct” removal shows a better performance, it still leaves outliers in the scene, negatively influencing the visual quality.
@@ IV-G1 NeRDS 360
Figure 6: SCRREAM scene01 (images cropped to POI).
TABLE III: Initialization iterations (1k-4k). We evaluate after 7k/30k final iterations on 6thAndMission_medium10 [14].
initial 7k final iterations 30k final iterations
@@
IV-G2 SCRREAM
To evaluate our approach in multiple settings, we also compared it on an indoor dataset. As shown in Figure 6, the same issues for GG as for the NeRDS dataset can usually be observed. The SCRREAM dataset [13] (real-world, indoor) consists of diverse scenes with highly accurate ground truth for semantic segmentation. As shown in Table IV, our approach outperforms GG in two scenes. In the second scene, a lower visual quality can be observed for both approaches, with GG showing a better result, in this case. Here, CoRe-GS is challenged and leaves in floaters as well, resulting in the low visual quality. Although this problem can be mitigated using a different $t_{r}$ value, we choose to keep the same hyperparameters for all scenes, to retain a fair comparison. However, a potential optimization can be seen in Fig. 7. Despite scene02, overall we outperform GG with an improvement in PSNR of 0.318, in SSIM of 0.046, and in LPIPS of 0.064.
Figure 7: Limitation. Exemplary image from scene02 of SCRREAM illustrating a failure case of CoRe-GS. The masked GT (left) is a rendering result from Table IV, and CoRe-GS using $t_{r}=1.0$ (right), instead of $t_{r}=0.5$ (center).
TABLE IV: Evaluation on SCRREAM [13]. We use the scenes 01/02/02 (_full_00) and select “chair” as POI.
$G$ (direct) [8] Ours
[…]
01 45.521 0.973 0.018 2614 45.799 0.994 0.014 709
02 21.663 0.829 0.056 2766 15.207 0.920 0.262 1083
03 45.521 0.973 0.018 2645 52.653 0.997 0.009 656
Mean 37.568 0.925 0.031 2675 37.886 0.970 0.095 816
@@ V DISCUSSION
We propose CoRe-GS, a novel method to refine POI for GS. Complete reconstruction of a scene is computationally intensive and often unnecessary when only certain POIs are relevant for timely decision-making. Our approach solves this problem with a hybrid solution: while a segmentation-ready scene can be examined early on, high-quality refinements are tailored to POIs, reducing the total training time to 25% compared to full semantic GS training. Our approach leverages semantic GS [8] in the first stage to initialize the POI refinement step. The semantic GS module is interchangeable and independent from the POI refinement, fostering modularity and scalability. Our color-based filtering and POI selection enable an efficient training cycle, ensuring that only data from the target object contributes to the scene optimization. This results in a more focused and faster training process by ignoring irrelevant background information.
As shown in Table I, the runtime of GG is notably higher compared to our approach. While semantic information is necessary for selecting POIs after obtaining the coarse, segmentation-ready scene representation, our method provides a substantial speedup and extracts more visually appealing POIs (Figure 3).
Our main motivation lies in situations where operators can select POIs from drone footage, which is why we focus mainly on outdoor data within the NeRDS 360 dataset. Experiments on the SCRREAM dataset validate the robustness of our method across diverse environments and highlight that further improvements may increase this robustness in indoor settings. Although the relative advantage over the baselines is smaller than on the outdoor NeRDS 360 scenes, our method still produces strong results. This translation to datasets outside of our intended application use case shows the scalability of our approach.
Additionally, the POI refinement gives CoRe-GS an advantage in occluded areas, see Fig. 3. In the complete scene, a pole or street light can occlude the POI from different NVS perspectives. Selecting a POI and extracting offers a more complete understanding of the object’s geometry and structure from the same perspective. This is not achievable with GG and “convex hull” removal. While direct removal allows it, outliers still negatively impact the result compared to our outlier-free visualization.
@@ VI LIMITATIONS
CoRe-GS benefits from high-quality segmentation. Its performance is therefore constrained by the segmentation quality. This limitation is not unique to our approach [4]. Previous work [4] has shown that the accuracy of segmentation quality has a critical influence onNVS quality and downstream tasks.
@@ VII CONCLUSION
We introduced CoRe-GS, a POI refinement for GS that selectively enhances segmentation-based POIs critical for scene exploration and operator decision making. By focusing computation on targeted scene areas, our approach produces higher-fidelity POIs without the overhead of refining the entire scene representation. Our evaluation on two benchmarks demonstrates clear advantages over both standard GS and GG in two object extraction settings. Unlike standard GS, which lacks POI support, our method achieves targeted visual refinement and, compared to GG, it reduces training time while preserving segmentation accuracy and rendering quality. These results highlight our approach’s efficiency and visual quality, while its modular design offers potential for future applications in more complex and specialized environments and mobile reconstruction scenarios.